OpenStudy (anonymous):

Four men and eight women are going to be assigned to a specific row of seats in the theater. if the 12 tickets for the numbered seats are given out at random, find the probability that four women are given the first four seats next to the center aisle.

OpenStudy (dan815):

@Mimi_x3

OpenStudy (mimi_x3):

lol i hate you dan Ok so here is what i propose There are 4 seats and 12 ppl; 4 men 8 women We wanna know what the probability that all 4 seats are women ( Lets forget abt the other seats cuz I dont think we even care abt them) Ok so the first seat the probability that it will be a woman is 8/12 Then the next seat .. one woman has been seated so there are 7 women and 11 ppl in total left SO the probability that the next seat will be a woman too is 7/11 Using the same logic the next seat will be 6/10 And obviously the last seat will be 5/9 SOOO the probability that all 4 seats will be women is : \[\frac{8}{12}*\frac{7}{11}*\frac{6}{10}*\frac{5}{9}=.14\]

OpenStudy (mimi_x3):

amirite boys?

OpenStudy (anonymous):

The answer was wrong

OpenStudy (anonymous):

as a fraction maybe?

OpenStudy (dan815):

ok show me your work so far

OpenStudy (mimi_x3):

dan was i right? -.-

OpenStudy (dan815):

? i just got here

OpenStudy (mimi_x3):

read my crap

OpenStudy (dan815):

heyam let me see your attemp so far

OpenStudy (anonymous):

the answer was right, i just needed to convert it to a fraction

OpenStudy (anonymous):

Thanks!

OpenStudy (mimi_x3):

MEDAL +FAN I dont work for free

OpenStudy (mimi_x3):

I needa make an income somehow or other